# S4ST: A Strong, Self-transferable, faSt, and Simple Scale Transformation for Data-free Transferable Targeted Attack

### v2

@@ 4.3 Comparison Against Existing Transformations
This subsection presents comparisons of the proposed S4ST with existing transformation methods across various settings.
@@ 4.3.1 Single Surrogate Attacks
Table IV reports the single surrogate model (ResNet-50) evaluation. Herein the transformed copies per iteration for SSA, SIA, BSR, and DeCoWa are limited to 1. We also set $m_{1}=1,m_{2}=5$ for Admix to ensure a fair comparison with SI. The results demonstrate that the proposed S4ST exhibits superior performance, not only achieving significantly higher average tSuc than existing transformation techniques, surpassing H-Aug by 14.2% and ODI by 24.9%, but also leading in terms of efficiency, with computational time only 25.7% of H-Aug and 28.4% of ODI. In fact, as can be seen in conjunction with Fig. 13, even the simple scaling transformation and our S4ST-Base are sufficient to achieve transferability on par with all methods except for T/H-Aug. And the further enhancements through block-wise scaling and complementary augmentations synergistically contribute to the final superior performance. This comparison is extended to alternative surrogate models (without further tuning the parameters), with results reported in Table V, and various attack iterations in Fig. 1. Our method consistently outperforms the others by clear margins.
TABLE V: Results by alternative surrogate models. strands for TTA-oriented methods. The best results are in bold.
Surrogate DenseNet-121 [69] VGG16-bn [70] ViT [60]
Attack tSuc Time tSuc Time tSuc Time
TMI [8, 9] 1.1 1.86 0.8 1.66 0.0 3.99
+ODI [32] 46.2 5.57 38.3 5.26 31.1 7.70
+SIA [11] 49.9 2.39 33.7 2.30 22.1 4.63
+DeCoWA [13] 43.6 3.89 46.5 3.66 14.6 8.30
+BSR [12] 47.9 2.03 32.1 1.95 27.6 4.27
+T-Aug [27] 49.3 2.01 48.1 1.75 27.2 4.06
+H-Aug [27] 59.3 4.51 48.3 7.00 - -
+S4ST (Ours) 74.6 2.15 65.6 1.83 64.8 4.05
@@ 4.3.2 Ensemble-based Attacks and Against Robust Victims
We further extend our comparison to scenarios involving the generation of AEs based on an ensemble of surrogate models and evaluations using robust victim models. For the selection of ensemble surrogate models, we distinguish between the commonly used self-ensemble in untargeted transfer attacks [46, 11, 13, 12], where the average gradient over multiple transformed copies is computed in a single iteration, and cross-model ensemble [8], which involves calculating the loss function using the average logits from different models to obtain gradients. In line with the evaluation protocols of previous research [15, 18], we incorporate robust victim models for evaluation, including Augmix [71], stylized training (SIN and SIN-IN) [72], adversarial training (AT) [2, 73], and Ensemble AT [3]. Table VI presents the results for tSuc and untargeted transfer success rate (uSuc) under the aforementioned settings.
TABLE VI: Results of attacking robust models (tSuc/uSuc) and ensemble-based attacks. stands for TTA-oriented methods. The best results are in bold. uSuc is the ratio (in %) of the false predictions caused by the targeted AE by the original correct predictions.
Attack Augmix [71] Stylized [72] Adversarial training (AT) [2] Ensemble AT [3] Avg. AvgC AvgT
[…]
@@
TABLE VII: Comparison with SOTA under the 10-Targets (all-source) setting. ##\## Data means the number of samples involved in method training. means the data is from the same dataset of victim models. means the victim models are involved for parameter tuning. The best results are in bold.
Attack Venue ##\## Data CNNs Vision transformers Robust mechanisms Avg.
@@
DFA [14] CVPR 2020 50K‡◆ 25.3 1.6 15.0 16.3 13.8 0.7 3.1 1.6 1.7 3.9 19.9 7.8 56.4 0.3 0.3 11.2
TTP [15] ICCV 2021 50K‡◆ 54.7 38.4 56.6 57.9 60.6 30.0 35.6 24.4 28.6 11.4 68.7 27.8 84.1 2.9 5.4 39.1
M3D [18] CVPR 2023 50K‡◆ 84.0 65.9 83.5 85.9 84.4 60.8 67.7 62.0 62.2 41.0 88.0 62.9 94.6 5.2 9.2 63.8
C-GSP [37] ECCV 2022 1.2M‡◆ 57.2 20.5 47.1 48.5 46.7 22.6 26.5 19.7 22.1 9.8 50.7 17.4 72.1 1.4 1.8 30.9
IAA [19] ICLR 2022 5K‡◆ 63.2 30.7 55.2 58.7 54.6 15.9 33.2 24.6 27.5 17.0 81.6 42.7 98.7 1.4 2.4 40.5
DRA-DI-TMI [20] IEEE TIP 2022 1.2M‡◆ 58.3 69.2 58.9 67.9 61.7 69.1 55.9 63.1 55.6 46.6 86.5 90.2 94.3 73.4 71.8 68.2
LRS-DI-TMI [35] NeurIPS 2021 1.2M‡◆ 72.3 55.1 61.3 72.9 60.5 45.6 47.2 50.9 41.5 35.1 85.0 41.5 93.3 24.5 31.7 54.6
SASD-WS-DI-TMI [21] CVPR 2024 1.2M‡◆ 88.6 85.2 92.2 93.3 91.2 58.2 70.4 68.8 61.6 47.2 95.3 88.3 99.8 12.4 20.8 71.6
SU-DI-TMI [33] CVPR 2023 N/A 54.7 17.1 41.6 43.0 33.2 6.5 19.2 18.8 11.7 10.2 43.8 11.2 89.1 0.8 1.0 26.8
CFM-RDI-TMI [29] CVPR 2023 N/A 88.1 81.7 89.0 91.0 88.2 56.3 72.3 72.7 65.3 49.4 96.0 79.6 99.7 17.0 26.5 71.5
ODI-TMI [32] CVPR 2022 N/A 74.3 62.6 73.6 75.1 70.8 37.6 53.3 55.5 42.8 32.2 78.1 45.2 96.8 6.1 9.6 54.2
H-Aug-TMI [27] ArXiv 2023 N/A 86.3 69.0 84.2 82.8 82.5 42.9 72.3 65.2 57.2 40.9 80.8 36.4 92.7 5.5 8.7 60.5
S4ST-TMI Ours ${}^{\quad\!\!\!}$ 96.0 86.2 94.5 95.0 94.7 69.1 88.0 84.9 80.1 66.5 95.8 68.1 99.5 18.1 28.3 77.7
Firstly, the proposed S4ST transformation significantly outperforms existing methods across all scenarios. Moreover, upon closer examination, it is evident that S4ST benefits the most from increased computational effort and a larger number of surrogate models—whether it is averaging over multiple variants of a single surrogate model or computing gradients across models. Notably, it achieves performance close to that of the cross-model ensemble in the self-ensemble case, highlighting its potential for practical application. This is particularly relevant since obtaining multiple distinct surrogate models is generally challenging in the black-box scenario. Additionally, adversarially trained models with larger perturbation budgets exhibit significantly improved defensive capabilities, yet they remain susceptible to erroneous (untargeted) predictions induced by targeted AEs.
4.4 Comparison with SOTA
We further compare with SOTA approaches for both data-free and -reliant TTAs. Note that training universal perturbations [14] or single-target generators [15, 18] for all target labels specified in the ImageNet-Compatible dataset is computationally impractical. Therefore, we follow the 10-Targets (all-source) setting [15] that averages tSuc on 10 target labels—{Grey-Owl, Goose, Bulldog, Hippopotamus, Cannon, Fire-Truck, Model-T, Parachute, Snowmobile, Street sign} against the entire dataset. All methods are implemented targeting ResNet-50 as the surrogate model and the attack results on three types of models—CNNs, Vision Transformers, and robustly trained models—are reported in Table VII.
The proposed method achieves the best result with an average tSuc of 77.7%, outperforming the current best data-free method, CFM [29], and the best data-reliant method, SASD-WS [21], by 6.2% and 6.1%, respectively, even when they are already equipped with input transformation techniques. Moreover, this is achieved in a strict black-box manner. As summarized in the table, SASD-WS utilizes 1.2M samples of the ImageNet-1K training set for post-training of the surrogate model and CFM also requires the 1K evaluated data and access to black-box models for parameter tuning. Overall, our method offers the flexibility to handle any (sample, label) pair while achieving SOTA targeted transferability to date. It does not utilize any black-box models or their data during the method design and training process, further revealing the potential and risks of TTAs under strict black-box conditions.
TABLE VIII: Target labels and semantically similar alternatives of attacking real-world APIs.
@@ 4.4 Comparison with SOTA
Baidu {CJK}UTF8gbsn指路牌, {CJK}UTF8gbsn指示牌, {CJK}UTF8gbsn路标 {CJK}UTF8gbsn鹅, {CJK}UTF8gbsn加拿大鹅, {CJK}UTF8gbsn肉鹅, {CJK}UTF8gbsn野鹅, {CJK}UTF8gbsn天鹅 {CJK}UTF8gbsn降落伞, {CJK}UTF8gbsn动力伞, {CJK}UTF8gbsn滑翔伞 {CJK}UTF8gbsn河马 {CJK}UTF8gbsn火炮
TABLE IX: Results of attacking real-world APIs. The best results are in bold.
Attack ##\## Data Google Cloud Vision Azure Baidu Avg.
@@
TABLE X: Results of attacking vision-language models. The best results are in bold.
Attack ##\## Data GLM-4V Flash [74] Gemini 1.5 Flash [75] Janus Pro 7B [76] Avg.
T1 T2 T3 T4 T5 Avg. T1 T2 T3 T4 T5 Avg. T1 T2 T3 T4 T5 Avg.
SASD-WS-DI-TMI [21] 1.2M 51 74 71 80 72 69.6 33 34 36 57 50 42.0 47 56 63 74 78 63.6 58.4
CFM-RDI-TMI [29] N/A 28 66 64 67 59 56.8 14 29 17 37 38 27.0 20 47 39 57 59 44.4 42.7
S4ST-TMI N/A 41 72 71 81 76 68.2 29 41 21 54 44 37.8 42 62 53 72 70 59.8 55.3
@@
4.5 Transferring Against Real-World API
Evaluation against the commercially deployed models in the real world is performed to verify the practicality of our method. We follow the quantitative evaluation setting of [18] to perform transfer attacks on 5 labels over 100 images against 3 commercial APIs including the Google Cloud Vision (label detection) [^5] https://cloud.google.com/vision/docs/labels, Azure (image tagging) [^6] https://portal.vision.cognitive.azure.com/demo/generic-image-tagging, and Buidu (advanced image recognition) [^7] https://cloud.baidu.com/product/imagerecognition/general. tSuc here is measured if the target label appears in the predicted label list of the APIs. Additionally, to prevent the scenario where the label set of the remote model does not include the target label, we first test the clean samples of the target label and select the high-frequency predicted labels returned by the APIs as semantically similar labels, as clarified in Table VIII.
Table IX presents the quantitative results. Our method achieves superior performance over the SOTA data-reliant approach SASD-WS [21] by 1.3% (69.6% average tSuc) when using a single ResNet-50 surrogate model while employing ensemble surrogates further elevates the average tSuc to 91.7%. Fig. 16 provides several examples obtained from these APIs, demonstrating that targeted AEs can induce high-confidence erroneous predictions toward the target class. These results demonstrate that our S4ST-empowered attacks hold significant potential not only for evaluating model vulnerabilities under unknown training data distributions and data processing pipelines but also for safeguarding against malicious exploitation of AI models in applications such as CAPTCHA systems.
4.6 Evaluation on Vision-language Models
To further validate the scalability of our method to vision-language models, we maintain the evaluation setting of the previous subsection and design a prompt to assess attack effectiveness in this open-vocabulary scenario. Specifically, we instruct the models to describe the main object(s) in the targeted AEs and then select the best match from six shuffled options, which include the source label, target label, and four randomly sampled labels from the ImageNet-1K label set.
Three popular efficient vision-language models are tested including GLM-4V Flash [74], Gemini 1.5 Flash [75], and Janus Pro 7B [76], with results shown in Table X. In this scenario, our method achieves an average tSuc of 55.3% across the three models, which, although 3.1% lower than SASD-WS method, still demonstrates its competitiveness in enabling effective transfer attacks without relying on data or training. Furthermore, by leveraging multiple surrogate models, the performance can be significantly improved to 82.4%. Fig. 17 illustrates the responses of vision-language models to several targeted AEs, showing that our method effectively induces target label-related visual features and corresponding textual understanding.
@@ 4.6 Evaluation on Vision-language Models
Figure 17: Responses obtained from vision-language models for targeted AEs generated by S4ST-TMI. Source labels are in green and target labels are in red.
4.7 Visual Inspections
Fig. 18 provides examples of targeted AEs. It demonstrates that our S4ST significantly enhances the infusion of the target label’s semantics into AEs, a process that is further augmented when attacking an ensemble of surrogate models. For instance, in the first set of examples, the bird crista is subtly altered to resemble the windows of a church loft; in the sixth set, the background among the flock of geese is imbued with the texture of a conch. These observations are consistent with previous findings on unbounded perturbations [24] and data-agnostic perturbations [15, 14], but they further underscore the manipulation of objects, textures, and backgrounds within individual images. More effective transformations can facilitate this process, revealing potential avenues for enhancing targeted transferability through transformations. However, this also reveals an intrinsic tension between stronger transferability and imperceptibility, which may provide a clue for distinguishing targeted AEs from untargeted ones and clean images, thereby aiding specializing defense strategies.

### v3

@@ 4.3 Comparison Against Existing Transformations
Competitors & Setup. We compare S4ST with DI [90], RDI [105], SI [40], Admix [77], SSA [47], ODI [2], SIA [78], BSR [76], T-Aug [83], and H-Aug [83]. For fair comparison, we limit transformed copies per iteration to 1 for SSA, SIA, BSR, and DeCoWa, and set $m_{1}=1,m_{2}=5$ for Admix. Unless otherwise specified, all gradient-based attacks adopt the same baseline approach described in Section 2.1, which incorporates TI [17], MI [16], and margin-based CE loss [84], with $T=900$.
@@ 4.3.1 Single Surrogate Attacks
Table III details the single-surrogate (ResNet-50) evaluation. S4ST exhibits dominant performance, surpassing H-Aug and ODI by 14.2% and 24.9% in average tSuc, respectively. It is also highly efficient, requiring only 25.7% and 28.4% of the computational time of H-Aug and ODI. Notably, aligning with Fig. 18, even simple scaling or S4ST-Base achieves transferability rivaling all baselines except T/H-Aug, while block-wise scaling and complementary augmentations synergistically secure the final SoTA performance. Furthermore, perturbations from all methods saturate under the $\ell_{\infty}$ constraint, exhibiting negligible $\ell_{2}$ norm differences. This consistent superiority extends to alternative surrogate models (Table IV, without retuning) and varying attack iterations (Fig. 3).
We also evaluate robustness against JPEG compression [23], a representative pre-processing defense. As shown in Fig. 20, RN50-generated AEs are subjected to varying JPEG qualities before victim model evaluation. Notably, S4ST maintains a high tSuc even when aggressive compression severely degrades the victims’ natural accuracy, demonstrating superior robustness. This resilience challenges the conventional generalization-robustness paradox, complicating defenders’ efforts to mitigate attacks without compromising legitimate model utility.
Figure 20: JPEG [23] defense evaluation. N/A indicates w/o defense and Q is the JPEG compression quality. AEs are generated by ResNet-50 and transferred against (a) CNX, (b) INv3, and (c) Swin.
@@ 4.3.2 Ensemble-based Attacks and Against Robust Victims
Table V presents the targeted (tSuc) and untargeted (uSuc) success rates for ensemble-based attacks against robust victims. We evaluate two strategies: self-ensemble [47, 78, 41, 76] (averaging gradients over multiple transformed copies per iteration) and cross-model ensemble [16] (averaging logits across different models). Following standard protocols [50, 99], our robust victim models include Augmix [28], stylized training [20], adversarial training (AT) [48, 56], and Ensemble AT [73].
S4ST significantly outperforms existing methods across all scenarios. Crucially, it scales exceptionally well with increased computational effort and additional surrogate models. In the self-ensemble setting, S4ST achieves performance rivaling the cross-model ensemble, a critical practical advantage, given the difficulty of obtaining diverse surrogate models in strict black-box scenarios. Furthermore, while adversarially trained models with larger perturbation budgets show stronger defenses, they remain highly susceptible to untargeted erroneous predictions induced by these targeted AEs.
4.4 Comparison with SoTA
Competitors & Setup. We benchmark S4ST-TMI against SoTA solutions: SASD-WS-DI-TMI [87], DRA-DI-TMI [103], IAA [104], Clean Feature Mixup (CFM-RDI-TMI) [3], Self-Universality (SU-DI-TMI) [82], Little Robust Surrogate (LRS-DI-TMI) [63], Dominant Feature Attack (DFA) [96], TTP [50], M3D [99], and C-GSP [92]. Because training universal perturbations [96] or single-target generators [50, 99] for all ImageNet-Compatible labels is computationally prohibitive (see supplementary), we adopt the 10-Targets (all-source) setting [50]. We average the tSuc on 10 target labels {Grey-Owl, Goose, Bulldog, Hippopotamus, Cannon, Fire-Truck, Model-T, Parachute, Snowmobile, Street sign} against the entire dataset. All methods target a ResNet-50 surrogate, evaluating across CNNs, ViTs, and robust models (Table VI).
TABLE V: Results of attacking robust models (tSuc/uSuc) and ensemble-based attacks. stands for TTA-oriented methods. uSuc is the ratio (in %) of the false predictions caused by the targeted AE by the original correct predictions.
  Attack Augmix [28] Stylized [20] Adversarial training (AT) [48] Ensemble AT [73] Avg. $_{\text{CNNs}} tSuc$$_{\text{ViTs}}$
SIN SIN-IN $\ell_{2}=.1$ $\ell_{2}=.5$ $\ell_{\infty}=.5$ $\ell_{\infty}=1.$ INv3ens3 INv3ens4 IRv2ens
@@ 4.4 Comparison with SoTA
TABLE VI: Comparison with SoTA under the 10-Targets (all-source) setting. #\# Data means the number of samples involved in method training. means the data is from the same dataset of victim models. means the victim models are involved for parameter tuning.
  Attack #\# Data CNNs ViTs Robust mechanisms Avg. $\ell_{2}$
CNX INv3 INv4 IRv2 Xcep ViT Swin PiT TNT DeiT Augmix SIN SIN-IN $_{\ell_{2}=.5} AT$$_{\ell_{\infty}=.5}$
@@
[…]
Quantitative Results. S4ST achieves a SoTA average tSuc of 77.7%, outperforming the best data-free (CFM [3]) and data-reliant (SASD-WS [87]) methods by 6.2% and 6.1%, respectively, despite them already utilizing input transformations. Crucially, S4ST operates strictly black-box. In contrast, SASD-WS post-trains surrogates using 1.2M ImageNet-1K samples, while CFM requires evaluated data and victim access for parameter tuning. Ultimately, our method offers unprecedented flexibility for any (sample, label) pair without relying on victims or external data, underscoring the severe risks of strict black-box TTAs.
Qualitative Results. Visualizations in Fig. 21 (and supplementary) demonstrate S4ST effectively infuses target semantics into AEs, an effect amplified by ensemble surrogates. For example, a bird’s crista morphs into church loft windows, and a goose flock’s background adopts conch textures. Consistent with findings on unbounded [100] and data-agnostic [50, 96] perturbations, this underscores the deliberate manipulation of local objects and textures. While robust transformations facilitate this semantic infusion, they inevitably expose an intrinsic tension between transferability and imperceptibility. This trade-off could inform specialized defenses by distinguishing targeted AEs from untargeted or clean images.
TABLE VII: Evaluation setting and results of attacking real-world APIs.
(a) Target labels and semantically similar alternatives
@@
Baidu 指路牌, 指示牌, 路标 鹅, 加拿大鹅, 肉鹅, 野鹅, 天鹅 降落伞, 动力伞, 滑翔伞 河马 火炮
(b) Quantitative results
  Attack #\# Data Google Cloud Vision Azure Baidu Avg.
@@
4.5 In-domain Task Transferability
We evaluate S4ST’s transferability across diverse natural image scenarios to validate its dataset- and task-agnostic nature. By directly transferring AEs crafted on an ImageNet-trained ResNet-50 to commercial APIs, VLMs, and dense prediction models without task-specific adaptation, we demonstrate S4ST captures universal semantic features that transcend varying architectures, training distributions, and learning paradigms.
4.5.1 Transferring Against Real-World API
Following [99], we attack Google Cloud Vision [^4] https://cloud.google.com/vision/docs/labels, Azure [^5] https://portal.vision.cognitive.azure.com/demo/generic-image-tagging, and Baidu [^6] https://cloud.baidu.com/product/imagerecognition/general APIs using 5 target labels across 100 images (500 AEs per API). An attack succeeds if the API’s returned labels include the exact target or pre-defined semantic equivalents (Table VII). As a negative control, we evaluate the success rate using clean images under the same protocol for Baidu; the baseline success rate is $0\%$, confirming that the observed tSuc is explicitly driven by our adversarial perturbations.
Using a single ResNet-50, S4ST achieves 69.6% average tSuc, outperforming the data-reliant SoTA SASD-WS [87] by 1.3%. Employing surrogate ensembles further elevates tSuc to 91.7%. Fig. 22 provides several examples obtained from these APIs, demonstrating that targeted AEs can induce high-confidence erroneous predictions toward the target class. These results demonstrate that our S4ST-empowered attacks hold significant potential for evaluating model vulnerabilities under unknown training distributions and data processing pipelines.
@@ 4.5.1 Transferring Against Real-World API
TABLE VIII: Results of attacking VLMs.
  Attack #\# Data GLM-4V Flash [30] Gemini 1.5 Flash [70] Janus Pro 7B [11] Qwen3-VL 8B [71] Avg.
T1 T2 T3 T4 T5 T1 T2 T3 T4 T5 T1 T2 T3 T4 T5 T1 T2 T3 T4 T5
SASD-WS-DI-TMI [87] 1.2M 51 74 71 80 72 33 34 36 57 50 47 56 63 74 78 43 78 69 86 75 61.4
CFM-RDI-TMI [3] N/A 28 66 64 67 59 14 29 17 37 38 20 47 39 57 59 21 65 49 79 65 46.0
S4ST-TMI N/A 41 72 71 81 76 29 41 21 54 44 42 62 53 72 70 43 73 62 85 72 58.2
@@
Figure 23: Responses obtained from VLMs for targeted AEs generated by S4ST-TMI. Source labels are in green and target labels are in red.
4.5.2 Evaluation on VLMs
To further validate the scalability of our method to VLMs, we maintain the evaluation setting used for commercial APIs and design a structured prompt to assess attack effectiveness in an open-vocabulary scenario. Specifically, for each AE, we provide the model with the following standardized instruction:
“You are an image analysis expert. Follow these steps strictly: 1. Description: Concisely describe the main object(s) in the picture in 1-2 sentences. 2. Selection: Choose ONLY ONE best match option from the list below marked with [ ]. Do NOT explain your choice. Options: {opt}. Output format: Separate description and selection with ’—’; Use ’[ ]’ to mark your final choice.”
The options $\{opt\}$ consist of six shuffled candidates: the source label, the target label, and four random ImageNet-1K labels. A successful attack forces the VLM to select the target label. As a negative control, we use the same prompt on clean images for Qwen3-VL 8B [71], where the baseline tSuc of selecting the target label is $0\%$ (accuracy of selecting the source label is $99\%$), confirming perturbations drive the deception.
Tested across GLM-4V Flash [30], Gemini 1.5 Flash [70], Janus Pro 7B [11], and Qwen-VL 8B [71] (Table VIII), S4ST achieves a competitive 58.2% average tSuc. While marginally lower (3.2%) than SASD-WS, S4ST rigorously maintains data-free constraints, demonstrating potent textual-visual manipulation without external data dependencies (Fig. 23).
4.5.3 Evaluation on Dense Prediction Tasks
TABLE IX: Results of transferring ImageNet surrogate (ResNet-50, target label: Model-T) agasint COCO2017 victims (corresponding label: car).
[…]
w/o attack 10.2 3.0 9.1 2.6 8.8 2.8
[…]
Instance segmentation demands precise spatial localization and pixel-level semantic discrimination. We transfer 1,000 AEs targeting Model-T (corresponding to car category of COCO [42]) to diverse architectures and detection paradigms trained on COCO: Mask R-CNN [26] (ResNet-101), Cascade Mask R-CNN [4] (CNX-S), and Mask2Former [12] (Swin-S) (Table IX). Metrics include $\text{tSuc}_{\text{Det}}$ (% of images detecting $\geq$1 car) and $\text{tSuc}_{\text{Seg}}$ (mean car mask area ratio).
S4ST exhibits remarkable cross-task transferability. On Mask2Former, it achieves 65.3% $\text{tSuc}_{\text{Det}}$ and 28.6% $\text{tSuc}_{\text{Seg}}$, significantly outperforming SASD-WS and CFM. Such high mask area ratios signify successful pixel-level contour manipulation (Fig. 24). Consistent gains across diverse backbones and detection heads confirm S4ST effectively exploits fundamental representation convergence within natural images, rather than merely overfitting to global semantics.
4.6 Cross-domain Adaptability (Boundary Cases)
To test our framework’s boundaries, we investigate cross-domain adaptability on multi-label thoracic disease classification and face verification. Unlike natural images, these domains feature fixed fields-of-view and standardized object scales. Evaluating them reveals whether our scaling-centric transformations and analytic measures generalize to structurally distinct data distributions. Detailed analyses are in the supplementary, with top-performing methods summarized in Fig. 25.
4.6.1 Multi-Label Thoracic Disease Classification
Evaluations regarding multi-label thoracic disease classification are conducted on the NIH CXR14 Dataset [79]. It contains over 100,000 chest X-ray (CXR) images, annotated with labels for 14 different thoracic conditions in a multi-label classification setting. We evaluate the Cardiomegaly label using 500 random test images (33 positive, 467 negative).
The surrogate model is an ImageNet-pretrained DenseNet-121 (DN121-IN). The victim models are {DN121-CXR, RN50-IN, ViT-S-CXR, ViT-B-CXR}, where -CXR denotes masked image modeling pretraining using CXR images [88]. This selection reflects variations in both training data distributions and model architecture. We generate AEs on negative images via Logit loss [100] and BIM [35] ($\epsilon=8/255,\alpha=1/255,T=100$), and then evaluate them alongside clean positive samples, and the AUC metric is computed for the black-box model. A lower AUC value indicates stronger transferability of the attack method.
For this task, the self-alignment analysis reveals that scaling transformations remain competitively effective. Meanwhile, the correlation analysis based on self-transferability yields conclusions consistent with those observed in natural image classification tasks. Therefore, we directly apply the aforementioned S4ST parameters without further modification. As shown in Fig. 25(a), S4ST significantly degrades the performance of classifiers with different pre-training data and network architectures, achieving superior performance.
Figure 24: Inference results by Mask2Former (Swin-S).
Figure 25: Visualization of targeted AEs and corresponding attack results on (a) NIH CXR14 thoracic disease classification and (b) CPLFW face verification.
4.6.2 Face Verification
We evaluate the face verification task using the CPLFW [101] dataset, which contains 12,000 aligned face images with diverse poses and is widely adopted in related works. We select 600 image pairs (300 positive, 300 negative) for evaluation.
The surrogate model is an Improved ResNet-101 trained with CurricularFace loss [32] on the MS1M [15] dataset (CurricularFace-IRN101-MS1M). The victim models consist of {FaceNet-IRv1-WebFace [58], FaceNet-IRv1-VGGFace2, ArcFace-IRN50-MS1M [15], and MobileFaceNet (MFN)-MS1M [15, 7]}. We attack one image per negative pair to maximize feature cosine similarity using BIM [35] ($\epsilon=16/255,\alpha=2/255,T=100$). The attacked negative pairs are then evaluated together with all positive clean pairs to obtain the ROC curves and corresponding AUC values of the victim models. Applying S4ST-Base ($p_{\textit{r}}=1.0,r=1.4,p_{\textit{aug}}=0,m=1$), our method demonstrates a clear performance gap, severely degrading victim AUCs (Fig. 25b).
Adaptation Strategy. In this case, the underlying task transitions from label-based classification to metric learning dominated by feature space distances, involving significantly different image distributions and data augmentation strategies [15]. Consequently, it becomes necessary to re-evaluate the overall transformation framework using our proposed self-alignment and self-transferability measures before conducting the BO search. Specifically, we observe a narrower tolerance for transformation intensity, along with strong correlations between geometric and color transformations. This likely stems from the limited scale diversity in fixed region of face images and the common practice of using minimal data augmentations. Our analysis reveals that directly applying the BO strategy for natural images (averaging self-transferability over 12 basic transformations) yields a weak PCC of 0.37 between the proxy and actual black-box transferability. However, by leveraging our self-transferability observations to simplify the framework, specifically by removing the redundant S4ST-Block and S4ST-Aug to focus solely on S4ST-Base, the PCC rises to 0.78. Further, by substitute the proxy by the self-transferability against scaling (findings provided in Fig. 11), the PCC reaches 0.97. This ensures that S4ST remains a principled and effective framework even under severe domain shifts. We also provide a streamlined adaptation strategy for such boundary cases. Please see supplementary.
5 Discussion: Why Scaling Excels?
In this section, we analyze why scaling transformations are distinctively effective in enhancing targeted transferability.
5.1 Why Transformations Benefit Transferability
As Section 3.2.1 establishes, the core function of transformations in gradient-based attacks is to restore the surrogate model’s feature distribution toward that of clean samples. But why do transformed samples maintain feature alignment?
The key is data variety and training-time augmentation. Neither CNNs nor ViTs possess inherent scale invariance; their transformation tolerance stems from the multi-scale nature of visual data and inductive biases learned via extensive augmentations. First, since natural images inherently contain objects at various scales, orientations, and perspectives, models must develop some degree of transformation tolerance to achieve generalizability. Second, training-time augmentations, exemplified by a scaling variation, RandomResizedCrop [^7] https://github.com/pytorch/vision/blob/main/torchvision/transforms/transforms.py#L855 (RRC), serve as a critical component across a wide spectrum of visual training paradigms, ranging from supervised to self-supervised learning (e.g., MAE [25], DINO [6]), from vision-only to vision-language tasks (e.g., CLIP [55]), and from classification to segmentation and detection, making it the de facto universal key to generalization capability in visual representation learning.
The shared data variety and dependency on a universal training component, designed to capture essential multi-scale data properties, rather than architectural priors, explains why scaling uniquely acts as a model-agnostic universal key for transferable AEs, a vulnerability S4ST explicitly exploits.
TABLE X: Accuracy of RN50 variants.
  Model RN50-stdRRC RN50-halfRRC RN50-woRRC
ImageNet val. set 76.1 75.2 69.0
ImageNet Compatible 93.2 90.6 79.3
5.2 Experimental Validation
Controlled experiments are conducted with three RN50 variants to validate our hypothesis:
  1. 1.
    RN50-stdRRC: official weights (v1) released by torchvision, trained with standard RRC augmentation (scale=(0.08,1.)).
  2. 2.
    RN50-halfRRC: the scaling intensity used is half that of the standard RRC (scale=(0.41,1.)).
  3. 3.
    RN50-woRRC: trained without RRC but the standard resize and center crop pipeline.
Their corresponding accuracies are reported in Table X. As expected, the model trained without RRC augmentation exhibits a significant drop in accuracy.
We evaluate the transferability of these models when used as surrogates and their robustness when attacked as victims. Fig. 26(a) shows that the effectiveness of transformation-based attacks diminishes as the intensity of RRC used in training surrogate decreases. Meanwhile, Fig. 26(b) exhibits that model robustness improves with weaker RRC intensity. These significant variance validates the crucial impact of RRC augmentation during training.
Figure 26: While stronger RRC augmentation enhances attack transferability and accuracy (a) at the cost of model robustness (b), models trained entirely without it still exhibit transformation tolerance (c) and the same correlation effects (d). This demonstrates that while training-time augmentation is a powerful modulator of adversarial properties, the fundamental basis is imparted by the intrinsic diversity of the data itself.
Notably, even without RRC (RN50-woRRC), scaling-related transformations like S4ST and H-Aug still yield the best transferability as attackers. Furthermore, when RN50-woRRC acts as a victim, it remains significantly more vulnerable to these scaling-based attacks than to other transformations. Additionally, Fig. 26(c) and (d) demonstrate that even without applying RRC during training, transformations can still help capture model alignment during attack, and strong internal correlations persist among geometric and color transformations. This indicates that the inherent diversity in the data, such as multi-scale, multi-orientation, and multi-perspective, plays a fundamental role in shaping the transformation-related adversarial properties.
5.3 Implications and Discussion
Why scaling excels. Our findings suggest that both the intrinsic diversity of the data itself and the use of training-time RRC augmentation collectively contribute to the efficacy of transformation-based TTAs, or more intrinsically, the transferability of AEs. And, the contribution of scale diversity appears to be significant than that of others. Scaling’s distinctive effectiveness originates from its direct correspondence to this data nature and coupled universal training practices established over decades of vision model development. This explains why prior attacks relying on arbitrary or partial aspects of data variety achieved limited success compared to S4ST.
The Dilemma of Generalization-Robustness Tradeoff. It also presents a dilemma analogous to that observed in adversarial training [48] and pre-processing defense [23], the generalization-robustness tradeoff. Specifically, models trained without RRC would likely suffer from compromised generalization capability while exhibiting stronger resistance to transfer attacks. Conversely, models trained with RRC achieve higher accuracy at the cost of increased vulnerability. This dual nature establishes data augmentation as a double-edged sword in visual representation learning. It connects adversarial vulnerability with broader robustness considerations and suggests that future secure model development must explicitly address this tradeoff.